OpenStudy (anonymous):

1. A weather balloon is launched from a height of 475 feet above sea level. If the balloon rises at a constant rate of 85 feet per minute, which equation could be used to determine t, the time in minutes it will take the balloon to reach a height of 9245 feet above sea level? A 9245 = 85 + 475t B 9245 = 85(t + 475) C 9245 = 475 + 85t D 9245 = (475 + 85)t 2. The perimeter of a rectangular wooden deck is 90 feet. The deck’s length, l, is 5 feet less than 4 times its width, w. Which system of linear equations can be used to determine the dimensions, in feet, of the wooden deck? Please

OpenStudy (anonymous):

@ShadowLegendX can you help me

OpenStudy (anonymous):

The perimeter of a rectangular wooden deck is 90 feet. The deck’s length, l, is 5 feet less than 4 times its width, w. Which system of linear equations can be used to determine the dimensions, in feet, of the wooden deck? * 2l + 2w = 90; l = 5 – 4w 2l + 2w = 90; l = 4 – 5w 2l + 2w = 90; l = 5w – 4 2l + 2w = 90; l = 4w – 5

OpenStudy (anonymous):

if you help i will give you a medal

OpenStudy (anonymous):

I need help for a huge test that will pass me in algebra

OpenStudy (shadowlegendx):

This is a test/?

OpenStudy (anonymous):

well part of one i got all the rest but these two i cant fighure out

OpenStudy (shadowlegendx):

We can't help you with tests, sorry :/

OpenStudy (anonymous):

its not a test its like a checkpoint it only has four answers on it and i ave to handwright it so its not really a test

OpenStudy (shadowlegendx):

Are you going to get a grade at the end of it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so can you help me

OpenStudy (anonymous):